OpenStudy (anonymous):

Find the next number in the number sequence 82, 101, 122, 145, ...

Directrix (directrix):

Suggested Answer -->50 --> Check this. The work came from the following link: http://answers.yahoo.com/question/index?qid=20101102185415AABydbI 145-122=23 122-101=21 101-82=19 82-65=17 65-50=15